OpenStudy (anonymous):

can someone please help me by finding the quadratic form of 9x^3/2 - x^3 - 8=0?

OpenStudy (mertsj):

Let a=x^3/2

OpenStudy (mertsj):

Then a^2=x^3

OpenStudy (radar):